OpenStudy (anonymous):

In the fall, you go to the local nursery and purchase a package of 10 rose plants. The clerk informs you that on average you can expect 5% of the plants to not live through the winter. Assume that the plants are a random sample. a) Use the binomial formula to determine the probability that one or more of the plants will not survive the winter. b) What is the probability that all the plants will survive? Write how the binomial formula is used to calculate the probability.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

a) what is the probability that 0 plants will not survive the winter?

jimthompson5910 (jim_thompson5910):

ie what is the probability that all plants survive?

OpenStudy (anonymous):

ummm how do i find that

jimthompson5910 (jim_thompson5910):

well the chances of one plant not making it (independent of the others) is 5% so the chances of it surviving is 95%

OpenStudy (anonymous):

Ok! so a is 95 percent?

jimthompson5910 (jim_thompson5910):

there are 10 plants, so the chances of all 10 plants making it is (0.95)^10 = 0.59873693923838 which is about a 59.87% chance

OpenStudy (anonymous):

Ohhh okay, i get that! that would be the answer for (a)

jimthompson5910 (jim_thompson5910):

now how can we use this to answer part a)

OpenStudy (anonymous):

you mean part b?

jimthompson5910 (jim_thompson5910):

think about it like this you have 2 possibilities a) all plants survive OR b) at least one (or more) dies

jimthompson5910 (jim_thompson5910):

no I haven't givin the answer to anything yet

jimthompson5910 (jim_thompson5910):

given*

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so if we know that the probability that all plants survive is 0.5987369 roughly and we have the two choices (given above) then the probability of at least one dying is what?

OpenStudy (anonymous):

um..

jimthompson5910 (jim_thompson5910):

remember all probabilities must add to 1

OpenStudy (anonymous):

41?

jimthompson5910 (jim_thompson5910):

think of a coin flip, you can either get heads or tails both have a chance of 0.5 so 0.5 + 0.5 = 1

jimthompson5910 (jim_thompson5910):

41 is way too big

OpenStudy (anonymous):

0.5987369 + x = 1?

jimthompson5910 (jim_thompson5910):

it need to be a number less than 1

jimthompson5910 (jim_thompson5910):

needs*

jimthompson5910 (jim_thompson5910):

yes solve that for x

OpenStudy (anonymous):

.4012631

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

and I just looked at part b again and I realized that I have answered it unintentionally

jimthompson5910 (jim_thompson5910):

part b is asking what is the probability that all plants survive and it's roughly 0.5987369

OpenStudy (anonymous):

Okay yah! what about a?

jimthompson5910 (jim_thompson5910):

what you got: 0.4012631

jimthompson5910 (jim_thompson5910):

basically P(all plants survive) + P(at least one dying) = 1 because those are the only two options

OpenStudy (anonymous):

oK:) so its : (a) (0.95)^10 = 0.59873693923 0.5987369 plus x = 1? x = 0.4012631 <----- (b) P(X = 10) = (0.95)^10 = 0.5987369 <----

jimthompson5910 (jim_thompson5910):

it's kinda like saying P(plant living) + P(plant dying) = 1 since a plant can only live or die

jimthompson5910 (jim_thompson5910):

good on both

OpenStudy (anonymous):

is the answer i put right

OpenStudy (anonymous):

ok thanks! can u look at my new problem? its just one more and its easier!!

jimthompson5910 (jim_thompson5910):

alright